how do u evaluate an expression like 2 to the 3rd power times 3 -5 times 5 to the 2nd power plus 8
Answer by roxy9229(32)

*=times
2^3*3-5*5^2

2*2*2=4*2=8 5*5=25

8*3-5*25+8
24-125+8

125-24=101
take sign of larger #: -101
-101+8=-93